If the length of a rectangular parking lot is 10 meters less than twice its width, and the perimeter is 400 meters, find the length of the parking lot.
Answer by ptaylor(2198)

Perimeter(p) of a rectangle=2*Length(l) + 2*Width(w)or p=2l+2w
Let w=width
Then length(l)=2*w-10
p=2(2w-10)+2w=400 get rid of parens (distributive law)
400=4w-20+2w add 20 to both sides
400+20=4w-20+20+2w collect like terms
420=6w divide both sides by 6
w=70 meters------------------------width
2w-10=2*70-10=140-10=130 meters -------length
CK
400=2*70+2*130
400=140+260
400=400
Hope this helps---ptaylor